# Asymmetric Effects Underlying Dynamic Heterogeneity in Miscible Blends of Poly(methyl methacrylate) with Poly(ethylene oxide)

### v1

@@ 2.2.1 Glass transition temperature
To obtain an apparent glass transition temperature ($T_{\text{g}}$), we use simulated dilatometry and analyze the temperature dependence of the specific volume, $v(T)$, for each system. An apparent $T_{\text{g}}$ is identified by a change in the slope of $v(T)$,65, 66 which reflects a shift in thermophysical relaxation behavior for the monitored quantity; this is associated with transition from a melt to a more glassy state. During the cooling phase of system preparation (Section 2.1.3), configurations are recorded at 10 K intervals. For each temperature, we perform an additional 1 ns simulation under isothermal–isobaric (NPT) conditions, starting from the corresponding configuration. The average specific volume is computed from the second half of each trajectory, yielding a set of $(T_{i},\bar{v}_{i})$ pairs.
The apparent $T_{\text{g}}$ from this dataset is determined using a previously reported bootstrap resampling procedure.20, 21 We first identify, by visual inspection, a maximum and minimum temperature within the melt regime, $T_{\text{max}}^{\text{m}}$ and $T_{\text{min}}^{\text{m}}$, and a maximum and minimum temperature within the glassy regime, $T_{\text{max}}^{\text{g}}$ and $T_{\text{min}}^{\text{g}}$. The temperatures $T_{\text{min}}^{\text{m}}$ and $T_{\text{max}}^{\text{g}}$ define sampling ranges for the melt, $[T_{\text{min}}^{\text{m}},T_{\text{max}}^{\text{m}}=T_{\text{min}}^{\text{m}}+100\leavevmode\nobreak\ \text{K}]$, and for the glass, $[T_{\text{min}}^{\text{g}}=T_{\text{max}}^{\text{g}}-100\leavevmode\nobreak\ \text{K},T_{\text{max}}^{\text{g}}]$. The temperature bounds used for each blend composition are listed in the Supplementary Information, Table S1. In each resampling iteration, a lower bound $T_{\text{lo}}^{\text{m}}$ is randomly selected from the melt range and an upper bound $T_{\text{hi}}^{\text{g}}$ from the glass range. Linear regressions are then performed on the data within $[T_{\text{lo}}^{\text{m}},T_{\text{max}}^{\text{m}}]$ and $[T_{\text{min}}^{\text{g}},T_{\text{hi}}^{\text{g}}]$ to fit the melt and glassy branches, respectively. The intersection of these fits defines a $T_{\text{g}}$ sample. This process is repeated 10,000 times to generate a distribution of $T_{\text{g}}$ values. We report the mean of this distribution as the estimated $T_{\text{g}}$ and its standard deviation as the associated uncertainty. Simulated $v(T)$ data and representative $T_{\text{g}}$ values for each blend composition are provided in the Supplementary Information, Figure S1.
@@ 2.2.2 Local segmental mobility
where $i$ is a backbone carbon, $\vec{r}_{i}(t)$ is the position of particle $i$ at time $t$, and $\langle\cdot\rangle_{\Delta t}$ denotes an ensemble average over the observation time. For the analysis herein, $\Delta t$ = 100 ps, which is substantially shorter than timescales for chain diffusion. Consequently, eq. (2) mostly captures local segmental fluctuations. This quantity is computed at 500 K, 360 K, and 220 K from the final 5 ns of the 35-nanosecond trajectories described in Section 2.1.3. Analysis based on $\mu_{i,\Delta t}$ across compositions and temperatures manifests in two ways. In Section 3.2, the segmental mobility is computed at the species level by averaging over all backbone carbons of each polymer type. In Section 3.3, the segmental mobility is further resolved based on the local environment of each backbone carbon atom to account for compositional heterogeneity introduced by blending.
To characterize the local environment of a particle, we define a normalized self-density parameter, $\tilde{\phi}^{(A)}_{i}$, which measures the local enrichment of species $A$ around a particle of the same species. For a given backbone carbon atom $i$, this quantity is computed as:
@@
$\omega_{ij}^{(A)}=\begin{cases}1&r_{ij}\leq\mu^{(A)}\\ \exp\left(-\dfrac{(r_{ij}-r_{m}^{(A)})^{2}}{\sigma^{2}}\right)&r_{ij}>\mu^{(A)}\end{cases}$ (4)
where $r_{ij}=|\vec{r}_{i}-\vec{R}_{j}|$ is the distance between particle $i$ and the center of mass of monomer $j$ (using the minimum image convention), and $r_{m}^{(A)}$ and $\sigma$ are parameters that define a smoothing kernel. To emphasize spatially local interactions, we set $r_{m}^{(A)}$ to the radius of gyration of a single monomer of species $A$, computed from simulations in vacuum at room temperature. This yields $r_{m}^{\text{(PEO)}}=1.59$ and $r_{m}^{\text{(PMMA)}}=2.34$. The smoothing width parameter is set to $\sigma=12$ based on the non-bonded, real-space interaction cutoff.
@@ 2.2.3 Free-volume analysis
The concept of free volume is often invoked to elucidate facets of polymer dynamics.67, 68, 69, 70 To quantify the free volume associated with each polymer chain, we implement the following procedure. First, the simulation cell is tessellated using Delaunay triangulation, such that each simplex (tetrahedron) is defined by four atoms. Each chain is associated with a subset of simplices that have at least one vertex belonging to an atom on that chain. Next, the entire simulation cell is filled with a three-dimensional grid of $n$ equally spaced spherical probes, where $n$ depends on a chosen probe radius. Each probe is then classified as occupied or unoccupied based on overlap with any atom in the system, using atomic diameters defined by the $\sigma$ parameters from the OPLS-AA force field. Finally, the free volume of a given chain is then computed as the total volume of unoccupied probes that fall within the chain-associated simplices. Free volume is computed and averaged over the final 5 ns of simulations equilibrated at 500 K, 360 K, and 220 K. In the main text, results correspond to a probe radius of 0.5 Å; additional results for other probe sizes are provided in Figure S2.
@@ 2.2.4 Rouse mode analysis
where $p$ = 0, $\dots$, $N-1$ indicates the mode index, $\vec{r}_{i}(t)$ is the position of the center of mass of the $i$th monomer on the chain at time $t$ , and $c_{p}$ is a $p$-dependent constant, such that $c_{1}=1$ and $c_{p}=2$ for all other modes. The zeroth mode corresponds to the behavior of the center of mass of a chain, while all other modes roughly correspond to the collective behavior of sub-chains of $\frac{(N-1)}{p}$ segments.
The Rouse mode coordinates are used to compute time autocorrelation functions (ACFs), which characterize the relaxation timescales of subchains of varying segment lengths. Each ACF is well-described by a stretched exponential function:
@@
$\tau_{p}^{\text{eff}}=\int_{0}^{\infty}\exp\left[-\left(\frac{t}{\tau_{p}}\right)^{\beta_{p}}\right]\text{d}t=\tau_{p}^{1/\beta_{p}}\Gamma\left(\frac{\beta_{p}+1}{\beta_{p}}\right)$ (8)
where $\Gamma(\cdot)$ denotes the gamma function. To perform the fitting, the ACF is computed over 8000 ps for PMMA and over 2000 ps for all other systems. These time intervals ensure that, at the highest temperature 500 K, the average normalized ACF of the largest $p$ mode of each species in each system reaches at least 0.5. For reference, representative ACFs and fitted curves are provided in the Supplementary Information, Figures S4 and S5.
@@ 3.1 Composition-dependent glass transition temperatures
Figure 1: Dependence of glass transition temperature ($T_{\text{g}}$) on blend composition. Markers represent simulated (filled) and experimental (empty) $T_{\text{g}}$ values for blends of PEO and PMMA. Experimental results are from Ref. 71. Solid black lines are fits to the Fox equation. Error bars reflect standard errors calculated from three independent system configurations. Horizontal, colored bands provide visual reference to 500 K (red), 360 K (purple), and 220 K (blue), which are examined in subsequent figures. The color gradation within each band distinguishes blends at the same temperature but different compositions.
We begin by briefly comparing experimental and simulated trends regarding $T_{\text{g}}$ for these PEO/PMMA systems as a function of composition. Figure 1 shows that both the simulated apparent $T_{\text{g}}$ as well as the experimental $T_{\text{g}}$ values71 follow the Fox equation
@@
where $w^{(i)}$ is the mass fraction of species $i$ and $T_{\text{g}}^{(i)}$ is the $T_{\text{g}}$ of a pure neat system of species $i$. The apparent $T_{\text{g}}$ values extracted from simulation are systematically higher than experimental values by approximately 100 K. This offset is roughly consistent across the range of examined compositions and is generally expected due to the much faster cooling rates and shorter observation times inherent to molecular simulations.72, 73 Given this systematic disparity, the similarity in trends and alignment with the Fox equation suggests that the employed force field captures the essential physics governing blend dynamics and responds appropriately to changes in composition.
Figure 1 also highlights three temperatures that are of specific interest in the following sections. These temperatures are selected to span distinct dynamical regimes: (i) 220 K lies below the apparent $T_{\text{g}}$ of both pure components and therefore below that of any blend; (ii) 360 K falls between the $T_{\text{g}}$ values of pure PEO and PMMA, such that some blends are above and others below their respective $T_{\text{g}}$; and (iii) 500 K exceeds the $T_{\text{g}}$ of both pure components and all blends. This temperature range enables the examination of how interspecies dynamical coupling depends not only on the different $T_{\text{g}}$ values of PEO and PMMA but also on the absolute temperature with respect to these $T_{\text{g}}$ values.
@@ 3.2 Characterization of species-dependent local dynamics
Figures LABEL:figMSFreltA and LABEL:figMSFreltB show the $\mu_{i,\Delta t}$ as a function of $T-T_{\text{g}}$, where $T_{\text{g}}$ varies with composition. The dynamics for PEO in blends (Figure LABEL:figMSFreltA) exhibit significant deviations from neat behavior at the same distance from $T_{\text{g}}$ (dashed black line) across most compositions but particularly in PMMA-rich blends (low $x^{\text{(PEO)}}$). This indicates that $T-T_{\text{g}}$ is not a reliable predictor of local segmental dynamics for PEO. In contrast, PMMA (Figure LABEL:figMSFreltB) dynamics in blends closely follow the behavior of neat PMMA, albeit with less strong correlation at temperatures below $T_{\text{g}}$. This trend is consistent with previous findings suggesting that PMMA dynamics are effectively governed by the temperature difference from the $T_{\text{g}}$ of the blend38, 39, 40 and also reveal an asymmetry in dynamical coupling between PEO and PMMA.
@@
To elucidate the prior results, we examine variations in species-dependent packing behavior, which is expected to manifest in different free volumes of the chains. Figures LABEL:figMSFreltC and LABEL:figMSFreltD show the normalized segmental mobility as a function of a change in free volume from the neat system (denoted by $\circ$’). The rationale for this comparison derives from considerations involving free-volume theory, which may suggest that $\log[\mu_{i,\Delta t}/\mu_{i,\Delta t}^{\circ}]\propto(\text{FV}-\text{FV}^{\circ})$. The data indeed possess roughly linear in the limit of smaller perturbations in FV. The positive trends in Figures LABEL:figMSFreltC and LABEL:figMSFreltD across all temperatures indicate that positive deviations in segmental mobility from neat polymer behavior generally correlate with increases in local free volume upon blending and vice versa. One exception is the dynamics of PMMA at high temperature, which is largely insensitive to changes in FV. Both PEO and PMMA chains in blends with lower $x^{\text{(PEO)}}$ exhibit greater free volume. We attribute this to less effective packing of PMMA relative to PEO due to sterics from the acrylate functional group. These trends support the notion that packing effects contribute to dynamic heterogeneity, but they again highlight an asymmetry in coupling. While increased free volume tends to correlate with enhanced mobility for PEO, a reduction in free volume does not universally imply suppressed segmental dynamics in PMMA by the same magnitude. This asymmetry reflects the influence of other factors beyond free volume in controlling relative enhancement/suppression of polymer dynamics in blends.

### v2

@@ 2.2.1 Glass transition temperature
To obtain an apparent glass transition temperature ($T_{\text{g}}$), we use simulated dilatometry and analyze the temperature dependence of the specific volume, $v(T)$, for each system. An apparent $T_{\text{g}}$ is identified by a change in the slope of v(T)v(T),74, 75 which reflects a shift in thermophysical relaxation behavior for the monitored quantity; this is associated with transition from a melt to a more glassy state. During the cooling phase of system preparation (Section 2.1.3), configurations are recorded at 10 K intervals. For each temperature, we perform an additional 1 ns simulation under isothermal–isobaric (NPT) conditions, starting from the corresponding configuration. The average specific volume is computed from the second half of each trajectory, yielding a set of $(T_{i},\bar{v}_{i})$ pairs.
The apparent $T_{\text{g}}$ from this dataset is determined using a previously reported bootstrap resampling procedure.20, 21 We first identify, by visual inspection, a maximum and minimum temperature within the melt regime, $T_{\text{max}}^{\text{m}}$ and $T_{\text{min}}^{\text{m}}$, and a maximum and minimum temperature within the glassy regime, $T_{\text{max}}^{\text{g}}$ and $T_{\text{min}}^{\text{g}}$. The temperatures $T_{\text{min}}^{\text{m}}$ and $T_{\text{max}}^{\text{g}}$ define sampling ranges for the melt, $T_{\text{min}}^{\text{m}},T_{\text{max}}^{\text{m}}=T_{\text{min}}^{\text{m}}+100\penalty 10000\ \text{K}]$, and for the glass, $T_{\text{min}}^{\text{g}}=T_{\text{max}}^{\text{g}}-100\penalty 10000\ \text{K},T_{\text{max}}^{\text{g}}]$. The temperature bounds used for each blend composition are listed in the Supplementary Information, Table S1. In each resampling iteration, a lower bound $T_{\text{lo}}^{\text{m}}$ is randomly selected from the melt range and an upper bound $T_{\text{hi}}^{\text{g}}$ from the glass range. Linear regressions are then performed on the data within $T_{\text{lo}}^{\text{m}},T_{\text{max}}^{\text{m}}]$ and $T_{\text{min}}^{\text{g}},T_{\text{hi}}^{\text{g}}]$ to fit the melt and glassy branches, respectively. The intersection of these fits is considered as one $T_{\text{g}}$ sample. This process is repeated 10,000 times to generate a distribution of $T_{\text{g}}$ values. We report the mean of this distribution as the estimated $T_{\text{g}}$ and its standard deviation as the associated uncertainty. Simulated $v(T)$ data and representative $T_{\text{g}}$ values for each blend composition are provided in the Supplementary Information, Figure S1.
@@ 2.2.2 Local segmental mobility
where $i$ is a backbone carbon, $\vec{r}_{i}(t)$ is the position of particle $i$ at time $t$, and $\langle\cdot\rangle_{\Delta t}$ denotes an ensemble average over the observation time. For the analysis herein, $\Delta t$ = 100 ps, which is substantially shorter than timescales for chain diffusion. Consequently, eq. (2) mostly captures local segmental fluctuations. This quantity is computed at 500 K, 360 K, and 220 K from the final 5 ns of the 35-nanosecond trajectories described in Section 2.1.3. Analysis based on $\mu_{i,\Delta t}$ across compositions and temperatures manifests in two ways. In Section 3.2, the segmental mobility is computed at the species level by averaging over all backbone carbons of each polymer type. In Section 3.3.1, the segmental mobility is further resolved based on the local environment of each backbone carbon atom to account for compositional heterogeneity introduced by blending.
To characterize the local environment of a particle in Section 3.3.1, we define a normalized self-density parameter, $\tilde{\phi}^{(A)}_{i}$, which measures the local enrichment of species $A$ around a particle of the same species. For a given backbone carbon atom $i$, this quantity is computed as:
@@
$\omega_{ij}^{(A)}=\begin{cases}1&r_{ij}\leq{\color[rgb]{0,0,0}r_{m}^{(A)}}\\ \exp\left(-\dfrac{(r_{ij}-r_{m}^{(A)})^{2}}{\sigma^{2}}\right)&r_{ij}>{\color[rgb]{0,0,0}r_{m}^{(A)}}\end{cases}$ (4)
where $r_{ij}=|\vec{r}_{i}-\vec{R}_{j}|$ is the distance between particle $i$ and the center of mass of monomer $j$ (using the minimum image convention), and $r_{m}^{(A)}$ and $\sigma$ are parameters that define a smoothing kernel. To emphasize spatially local interactions, we set $r_{m}^{(A)}$ to the radius of gyration of a single constitutional repeat unit of species $A$, computed from simulations in vacuum at room temperature. This yields $r_{m}^{\text{(PEO)}}=1.59$ Å and $r_{m}^{\text{(PMMA)}}=2.34$ Å. The smoothing width parameter is set to $\sigma=12$ Å based on the non-bonded, real-space interaction cutoff.
@@ 2.2.3 Free-volume analysis
The concept of free volume is often invoked to elucidate facets of polymer dynamics.76, 77, 78, 79 To quantify the free volume associated with each polymer chain, we implement the following procedure. First, the simulation cell is tessellated using Delaunay triangulation, such that each simplex (tetrahedron) is defined by four atoms. Each chain is associated with a subset of simplices that have at least one vertex belonging to an atom on that chain. Next, the entire simulation cell is filled with a three-dimensional grid of $n$ equally spaced spherical probes, where $n$ depends on a chosen probe radius. Each probe is then classified as occupied or unoccupied based on overlap with any atom in the system, using atomic diameters defined by the $\sigma$ parameters from the OPLS-AA force field. Finally, the free volume of a given chain is then computed as the total volume of unoccupied probes that fall within the chain-associated simplices. Free volume is computed and averaged over the final 5 ns of simulations equilibrated at 500 K, 360 K, and 220 K. In the main text, results correspond to a probe radius of 0.5 Å; additional results for other probe sizes are provided in Figure S2.
2.2.4 Distribution of local composition
To further probe how packing influences polymer dynamics, we perform an analysis akin to that described by the Lodge–McLeish (LM) chain connectivity model, which assumes that a cooperative volume of spanned by a Kuhn length $\ell_{k}$ governs local self-concentration effects.48 Here, we go beyond mean-field average self-concentrations and calculate the distribution of local compositions surrounding PEO and PMMA backbone carbons within spheres of radius equal to the Kuhn length, $\ell_{k}$.80 For PEO, we use $\ell_{k}$ = 8.2 Å; for PMMA, we use $\ell_{k}$ = 13.8 Å.73
The calculation proceeds similarly to that described in Section 2.2.3. First, the simulation cell is filled $n$ spherical probes distributed on a simple cubic lattice; $n$ depends on the probe radius, which is 0.5 Å for our analysis. Each probe is subsequently classified as PEO, PMMA, or unoccupied based on whether the center-to-center distance between a probe and a particle is less than the atomic diameter; atomic diameters are based on the values of $\sigma$ as prescribed by the employed force field; probes are also assigned to a given polymer chain based on which atom overlaps with the probe.
For each backbone carbon of polymer type $A$, we then quantify three local composition measures within the $\ell_{k}$ sphere: the intramolecular volume fraction $\phi_{\text{intra}}^{\text{(A)}}$ (i.e., contributions from the same chain), the intermolecular PEO volume fraction $\phi_{\text{inter,PEO}}^{\text{(A)}}$ (i.e., contributions from surrounding PEO chains), and the intermolecular PMMA volume fraction $\phi_{\text{inter,PMMA}}^{\text{(A)}}$ (i.e., contributions from surrounding PMMA chains). These quantities are calculated as the sum of volumes of appropriately classified probes divided by the total sphere volume. The total sphere volume is defined as the sum of intramolecular, intermolecular PEO, intermolecular PMMA, and free volume contributions within the relevant sphere. This normalization is slightly different from prior analyses of a similar nature, where normalization was done on a species-specific basis.80 Distributions of volume fractions are computed and averaged over the final 5 ns of trajectories equilibrated at 500 K, 360 K, and 220 K.
To further characterize these distributions, we calculate their Shannon entropy:
$H(x^{\text{(PEO)}};\phi_{\kappa}^{\text{(A)}})=\int P(\phi_{\kappa}^{\text{(A)}})\log_{2}\left[P(\phi_{\kappa}^{\text{(A)}})\right]d\phi_{\kappa}^{\text{(A)}}$ (5)
[…]
@@ 2.2.5 Rouse mode analysis
where $p$ = 0, $\dots$, N1N-1 indicates the mode index, $\vec{r}_{i}(t)$ is the position of the center of mass of the $i$th monomer on the chain at time $t$ , and $c_{p}$ is a $p$-dependent constant, such that $c_{1}=1$ and $c_{p}=2$ for all other modes. The zeroth mode corresponds to the behavior of the center of mass of a chain, while all other modes roughly correspond to the collective behavior of sub-chains of $\frac{(N-1)}{p}$ segments.
The Rouse mode coordinates are used to compute time autocorrelation functions (ACFs), which characterize the relaxation timescales of subchains of varying segment lengths. Each ACF is well-described by a stretched exponential function:40, 81
@@
$\tau_{p}^{\text{eff}}=\int_{0}^{\infty}\exp\left[-\left(\frac{t}{\tau_{p}}\right)^{\beta_{p}}\right]dt=\tau_{p}^{1/\beta_{p}}\Gamma\left(\frac{\beta_{p}+1}{\beta_{p}}\right)$ (9)
where $\Gamma(\cdot)$ denotes the gamma function. To perform the fitting, the ACFs were computed over time intervals of 8 ns for PEO, while for PMMA, intervals were 125 ns, 75 ns, and 20 ns for $x^{\text{(PEO)}}=$ 0.0, $x^{\text{(PEO)}}=0.1$–0.2, and $x^{\text{(PEO)}}=0.9$–0.3, respectively. These intervals were selected to ensure that the average normalized ACF of the largest Rouse mode for each species decayed to at least 0.2 at 500 K, to facilitate reliable fitting. Representative ACFs and fitted curves are shown in the Supplementary Information (S10 and S11).
@@ 3.1 Composition-dependent glass transition temperatures
Figure 1: Dependence of apparent glass transition temperature ($T_{\text{g}}$) on blend composition. Markers represent simulated (filled) and experimental (empty) $T_{\text{g}}$ values for blends of PEO and PMMA. Experimental results are from Ref. 82. Solid black lines are fits to the Fox equation. Error bars reflect standard errors calculated from three independent system configurations. Horizontal, colored bands provide visual reference to 500 K (red), 360 K (purple), and 220 K (blue), which are examined in subsequent figures. The color gradation within each band distinguishes blends at the same temperature but different compositions.
We begin by briefly comparing experimental and simulated trends regarding $T_{\text{g}}$ for these PEO/PMMA systems as a function of composition. Figure 1 shows that both the simulated apparent $T_{\text{g}}$ as well as the experimental $T_{\text{g}}$ values82 follow the Fox equation
@@
where $w^{(i)}$ is the mass fraction of species $i$ and $T_{\text{g}}^{(i)}$ is the $T_{\text{g}}$ of a pure neat system of species $i$. The PMMA used in the experimental measurements of $T_{\text{g}}$ are of unspecified tacticity.82 The apparent $T_{\text{g}}$ values extracted from simulation are systematically higher than experimental values by approximately 100 K. This offset is roughly consistent across the range of examined compositions and is generally expected due to the much faster cooling rates and shorter observation times inherent to molecular simulations.83, 84 Given this systematic disparity, the similarity in trends and alignment with the Fox equation suggests that the employed force field captures the essential physics governing blend dynamics and responds appropriately to changes in composition.
Experiments using differential scanning calorimetry (DSC) on PEO/PMMA blends have reported both single82, 85, 86 and dual87, 27 $T_{\text{g}}$ signals at intermediate compositions, with the former attributed to broad, overlapping peaks arising from local composition fluctuations.87 In general, rapid quench rates and finite system sizes inherent to simulations may obscure multiple signatures observed experimentally. Our simulations yield a single apparent $T_{\text{g}}$ based on the systematic fitting procedure described in the Supplementary Information (Section S1) However, recent simulation work28 suggested two $T_{\text{g}}$ values from density–temperature data using hyperbolic fitting. Our data (Figure S1) illustrate a broad transition window at intermediate compositions than could support hyperbolic fitting with two distinct inflection points. Regardless of whether the simulated dilatometry yields one or two apparent $T_{\text{g}}$, the key outcome is that the simulations produce a reasonable composition dependence, suggesting it can capture asymmetries in PEO and PMMA dynamics.
Figure 1 also highlights three temperatures that are of specific interest in the following sections. These temperatures are selected to span distinct thermal regimes: (i) 220 K lies below the apparent $T_{\text{g}}$ of both pure components and therefore below that of any blend; (ii) 360 K falls between the $T_{\text{g}}$ values of pure PEO and PMMA, such that some blends are above and others below their respective $T_{\text{g}}$; and (iii) 500 K exceeds the $T_{\text{g}}$ of both pure components and all blends. This temperature range enables the examination of how interspecies dynamical coupling depends not only on the different $T_{\text{g}}$ values of PEO and PMMA but also on the absolute temperature with respect to these $T_{\text{g}}$ values.
@@ 3.2 Characterization of species-dependent local dynamics
Figures LABEL:figMSFreltA and LABEL:figMSFreltB show the $\mu_{i,\Delta t}$ as a function of $T-T_{\text{g}}$, where $T_{\text{g}}$ varies with composition. The dynamics for PEO in blends (Figure LABEL:figMSFreltA) exhibit significant deviations from neat behavior at the same distance from $T_{\text{g}}$ (dashed black line) across most compositions but particularly in PMMA-rich blends (low $x^{\text{(PEO)}}$). This indicates that $T-T_{\text{g}}$ is not a reliable predictor of local segmental dynamics for PEO. In contrast, PMMA (Figure LABEL:figMSFreltB) dynamics in blends closely follow the behavior of neat PMMA, albeit with less strong correlation at temperatures below $T_{\text{g}}$. This trend is consistent with previous findings suggesting that PMMA dynamics are effectively governed by the temperature difference from the $T_{\text{g}}$ of the blend41, 42, 43 and also reveal an asymmetry in dynamical coupling between PEO and PMMA. The mobilities of both PEO and PMMA at 220 K (below $T_{\text{g}}$ of both pure components) decrease with increasing $x^{\text{(PEO)}}$. This behavior is both opposite to the behavior at 360 K and 500 K and also possibly unexpected because PEO is the higher-mobility species in the blend.
@@
To elucidate the prior results, we examine variations in species-dependent packing behavior, which is expected to manifest in different free volumes of the chains. Figures LABEL:figMSFreltC and LABEL:figMSFreltD show the normalized segmental mobility as a function of a change in free volume (FV) from the neat system (denoted by $\circ$’). The rationale for this comparison derives from considerations involving free-volume theory, which may suggest that $\log[\mu_{i,\Delta t}/\mu_{i,\Delta t}^{\circ}]\propto(\text{FV}-\text{FV}^{\circ})$. The data indeed possess roughly linear behaviors in the limit of smaller perturbations in FV. The positive trends in Figures LABEL:figMSFreltC and LABEL:figMSFreltD across all temperatures indicate that positive deviations in segmental mobility from neat polymer behavior generally correlate with increases in local free volume upon blending and vice versa.
Increases in FV also explain the apparent enhancement in mobility of both PEO and PMMA at 220 K with increasing mole fraction of the lower mobility species. In particular, both species are effectively glassy at 220 K, with low mobilities of similar magnitude ($\mu^{\circ}_{\text{PEO},\Delta t}=0.16$ Å/ns and $\mu^{\circ}_{\text{PMMA},\Delta t}=0.09$ Å/ns), such that the relative mobility advantage of PEO is diminished (Figure S4). Consequently, in this regime, we suggest that the dynamics are mostly controlled by variations in free volume. Although PMMA is canonically slower, it is also bulkier, such that its addition leads to increases FV, and thus less effective packing. By contrast, PEO-rich blends exhibit lower mobilities due to more efficient packing of chains. The less effective packing of PMMA relative to PEO is supported by the observation that the free volume of neat PMMA surpasses that of neat PEO at the same $T-T_{\text{g}}$ (Figure S3).
An exception to the observation that dynamics are controlled by free volume relates to PMMA at high temperatures. In this case, mobility in blends resembles that of a $T_{\text{g}}$-equivalent neat polymer reference, despite variations in FV (Figure LABEL:figMSFreltD). These trends support the notion that packing effects contribute to dynamic heterogeneity, but they again highlight an asymmetry in coupling. While increased free volume tends to correlate with enhanced mobility for PEO, a reduction in free volume does not universally imply suppressed segmental dynamics in PMMA by the same magnitude. This asymmetry reflects the influence of other factors beyond free volume in controlling relative enhancement/suppression of polymer dynamics in blends.